# CLARITY: Contextual Linguistic Adaptation and Accent Retrieval for Dual-Bias Mitigation in Text-to-Speech Generation

### v1

@@ Abstract
Instruction-guided text-to-speech (TTS) research has reached a maturity level where excellent speech generation quality is possible on demand, yet two coupled biases persist: accent bias, where models default to dominant phonetic patterns, and linguistic bias, where dialect-specific lexical and cultural cues are ignored. These biases are interdependent, as authentic accent generation requires both accent fidelity and localized text. We present CLARITY (Contextual Linguistic Adaptation and Retrieval for Inclusive TTS sYnthesis), a backbone-agnostic framework that addresses these biases through dual-signal optimization: (i) contextual linguistic adaptation that localizes input text to the target dialect, and (ii) retrieval-augmented accent prompting (RAAP) that supplies accent-consistent speech prompts. Across twelve English accents, CLARITY improves accent accuracy and fairness while maintaining strong perceptual quality [^1] Code and audio samples are available at https://github.com/ICT-SIT/CLARITY.
Index Terms—  Instruction TTS, Accent Bias, Retrieval
1 Introduction
@@
Instruction-guided TTS systems [du2024cosyvoice, lyth2024natural, zhou2024voxinstruct, wang2025spark] taking two inputs: 1) transcript text and 2) speaker descriptions of desired speech characteristics, have achieved promising instruction-following abilities. However, when prompted with natural language instructions, such TTS systems are susceptible to various biases. While gender bias has been examined by prior work [kuan-lee-2025-gender], to the best of our knowledge, no study has jointly addressed accent and linguistic bias. The two are interrelated but accent bias stems from the TTS system, while linguistic bias is often implicit in user input. The dual bias, as illustrated in Fig. 1, reduces generated speech authenticity by misaligning what is said and how it is spoken [burns2019speaking]. For example, instructing TTS to speak in a British Accent but with American English oriented text [^2] Note that traditional TTS assumes fixed text. This work addresses scenarios where users want to generate speech in a specific accent but cannot provide text in that variant of English. Hence, the input text must be slightly adapted to the accent specified in the speaker instruction..
@@
Fig. 1: The “dual-bias” linguistic mismatch (prompt-source) with accent bias (system-source) reduces naturalness and authenticity.
@@
Linguistic Bias – in the context of TTS, this is generally not an issue of model failure but a human-centric problem rooted in users’ ingrained linguistic knowledge, shaped by education and culture [clark2021impact]. It often leads to standardized input text, even when dialectal diversity is intended, preventing authentic synthesis. Recent studies confirm this tendency: for instance, ChatGPT frequently defaults to “standard” English when prompted with non-standard dialect text [fleisig-etal-2024-linguistic].
@@
Accent Bias – arises from training data imbalance, where mainstream accents (e.g., American English) dominate and less common ones are under-represented. As a result, models may default to generic accents even when instructed otherwise. Recent research confirms performance disparities across accents in synthetic speech. For example, [10.1145/3715275.3732018] found significant quality gaps when cloning voices with different English accents, potentially reinforcing accent-based discrimination in AI speech. Broader evaluations like AudioTrust [li2025audiotrust] revealed that accent is a stronger source of unfair bias in audio models than age or gender, Current models generally default to American-accented speech profiles.
@@
CLARITY: To jointly mitigate this dual bias, we present Contextual Linguistic Adaptation and Retrieval for Inclusive TTS sYnthesis (CLARITY). Our proposed CLARITY is a backbone-agnostic framework to enhance zero-shot TTS by providing richer guidance for both linguistic content and accent style. Its novelty lies in a two-signal optimization to jointly mitigate the two biases through: (i) an LLM-guided text adaptation module to localize input to the target dialect, and (ii) a metadata-driven retrieval mechanism to supply accent-consistent prompts.
Our Contributions. We introduce a sociolinguistic innovation in TTS by using large language models (LLMs) for on-the-fly text style transfer, leveraging their knowledge of dialects and slang to enhance authenticity. By pairing this with retrieval-augmented prompting, CLARITY achieves accent control and dialect-aware synthesis. To our knowledge, this is the first framework to jointly address user-side linguistic bias and system-side accent bias. Objective and subjective evaluations show that CLARITY improves fairness, accent fidelity, and overall authenticity.
2 The Proposed CLARITY
@@
The CLARITY framework promotes inclusive TTS by mitigating linguistic and accent biases through two signals: an adapted text $x$ and an accent-consistent prompt $s$, as shown in Fig. 2. We cast this as a two-signal optimization problem, aiming to maximize both linguistic and accent fidelity under the synthesis constraint of a backbone TTS model. Formally, the objective is defined as
@@
where $J_{\text{LLM}}(x^{*},m)$ denotes the judgment score assigned by an LLM-as-a-Judge evaluating whether the adapted text $x^{*}$ aligns with the requested metadata $m$, and $C(s^{*},m)$ is the accent recognition confidence score that the retrieved prompt $s^{*}$ matches the target accent attributes. The synthesis function $g_{\text{TTS}}$ represents any zero-shot backbone TTS model.
LLM-Guided Instruction Parsing: Let $u\in\mathcal{U}$ be a free-form user instruction in natural language, and $m\in\mathcal{M}$ the structured metadata schema, $m=(\text{accent}\in\mathcal{A},;\text{gender}\in{M,F},;\text{age}\in\mathbb{N},\ldots)$, where $\mathcal{A}$ is the set of supported accents. The challenge is that $u$ may encode attributes explicitly, vaguely, or implicitly; for example, instead of the prescriptive instructions shown in Fig. 2, there may be one open to interpretation such as u=u=Read this like a local in Singapore” (a locale with multiple races and accents).
We employ an LLM to parse $u$ into the structured metadata $m$, with the following parsing function:
@@
where $f_{\text{parse}}$ is realized by prompting the LLM with schema-specific instructions, and $\theta_{\text{LLM}}$ are its parameters. For each attribute $m_{j}$ (e.g., accent, gender, age), the LLM estimates a posterior distribution, $P(m_{j}\mid u;\theta_{\text{LLM}}),\quad m_{j}\in\mathcal{V}_{j},$ where $\mathcal{V}_{j}$ denotes the vocabulary/domain of attribute $j$. The predicted value is chosen as $m_{j}=\arg\max_{v\in\mathcal{V}_{j}}P(m_{j}=v\mid u;\theta_{\text{LLM}}).$ In cases where $u$ provides insufficient information to infer slot $m_{j}$, the system falls back to default $\pi_{j}$.
Contextual Linguistic Text Adaptation: Let $x$ denote the input text provided by the user, linguistic bias on the user side arises because $x$ is often drawn from the distribution of the user’s learned dialect $\mathcal{D}_{u}$, which may differ from the target dialect $\mathcal{D}_{t}$. We formalize text adaptation as a conditional function as follows:
@@
To ensure that the adapted text $x^{\prime}$ is linguistically faithful to the target dialect $\mathcal{D}_{t}$, we generate candidate adaptations $\{x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{K}\}$ using different LLMs (e.g., GPT and LLaMA). We compare three candidates: the original standard text $x_{\text{std}}$ and adapted texts $k\in\mathcal{K}=\{\text{GPT},\text{LLaMA}\}$ produced by different LLMs. An LLM-as-a-Judge assigns each candidate a score $J_{\text{LLM}}(x,m)\in[0,10],$ which reflects how well $x$ aligns with the metadata $m$, particularly the accent attribute. The final text is then chosen as the candidate with the highest judgment score:
@@
Fig. 2: Our proposed CLARITY for mitigating user-side linguistic bias and system-side accent bias.
Retrieval-Augmented Accent Prompting (RAAP): System-side accent bias arises when a TTS backbone trained on imbalanced data defaults to majority accents $\mathcal{A}_{\text{dom}}$. To counter this, we introduce Retrieval-Augmented Accent Prompting (RAAP), which selects accent-consistent prompts from a curated pool $\mathcal{P}=\{(a_{i},z_{i},s_{i})\}$ containing metadata, transcript, and speech. Given target metadata $m$, we filter $\mathcal{P}$ by $a_{i}\approx m$ to form $\mathcal{P}_{m}$, then score each candidate $s_{i}$ using an ECAPA-TDNN [desplanques2020ecapa] for accent confidence $C(s_{i},m)\in[0,1]$, and select those with the highest scores to guide accented speech generation. Mathematically,
@@
To further align the text content of the prompt speech with the user’s text input, RAAP calculates the text similarity between the prompt speech candidates and the user-provided (standard) text input using $\cos(\phi(z_{i}),\phi(x))$, where $\phi(\cdot)$ is a TF–IDF [salton1988term] text embedding representation and $\cos(\cdot,\cdot)$ denotes cosine similarity. Specifically, we define the effective ranking score as
@@
Backbone-agnostic Guided Synthesis: The final stage of CLARITY aligns the two signals: the adapted text $x$ and the accent-consistent prompt $s$, for guided synthesis. Let $g_{\text{TTS}}$ be the backbone TTS with parameters $\theta_{\text{TTS}}$. As a backbone-agnostic framework, CLARITY can use any instruction-guided, zero-shot TTS model; here we adopt the SOTA CosyVoice2 [du2024cosyvoice2] with its zero-shot functionality. The synthesis is defined as
@@
3 Experiments
3.1 Experimental Settings
@@
Accent Pool Construction: We evaluated our method on twelve English accents: ten accents (Canadian (CA), Chinese (CN), Spanish (ES), British (GB), Indian (IN), Japanese (JP), Korean (KR), Portuguese (PT), Russian (RU), and American (US)) are from the AESRC dataset (52,614 utterances from 528 speakers) [shi2021accented], each contributing 44–92 speakers and 4.2k–7k utterances, balanced by gender and ages 15–69. The remaining two accents are drawn from the SEAME dataset [lyu2010seame]: 40 Malaysian speakers (4,358 utterances, aged 20 to 33) and 114 Singaporean speakers (17,750 utterances, aged 18 to 24), with code-switching between English and Mandarin.
@@
Text and Free-form Instructions Curation: For transcript, GPT-4333https://openai.com/index/gpt-4-research/ generated standard sentences across four scenarios (restaurant, university, workplace, supermarket). Speaker instructions derived from accent-pool metadata (accent, age, gender) served both as prompts and ground-truth labels for retrieval. Each instruction–text pair was then adapted by GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ and LLaMA-3.1-8B [^5] https://huggingface.co/meta-llama/Llama-3.1-8B to produce accent-specific variants. Fidelity was evaluated by GPT-5666https://openai.com/index/introducing-gpt-5/, which rated text–accent alignment on a 1–10 scale.
@@
Baseline Systems: Two open-source SOTA prompt-based TTS systems are selected as baselines: ParlerTTS [lacombe-etal-2024-parler-tts, lyth2024natural] and CosyVoice2 [du2024cosyvoice]. Both take user instructions and standard text as input, but CosyVoice2 requires additional prompt speech even for instruction-guided functionality [^7] https://github.com/Render-AI/CosyVoice2 where a silent speech sample is provided.
@@
Object Evaluation Metrics: To quantify the accent level of the generated speech, we calculate accent accuracy using the ECAPA-TDNN accent recognition model [^8] https://huggingface.co/Jzuluaga/accent-id-commonaccent_ecapa, finetuned on the accent pool (Section 3.1), as in RAAP. Speech quality is measured with NISQA [mittag2021nisqa] (1–5 scale) using its pre-trained model [^9] https://github.com/gabrielmittag/NISQA. Bias/Fairness of the generated speech is evaluated with the Fairness Discrepancy Rate (FDR) [de2021fairness, estevez2023study], which quantifies disparities in false alarm/reject rates across accents, with 1 indicating perfect fairness.
@@
Subject Evaluation Metrics: We conducted human listening tests on four accented English samples (CN, SG, IN, GB). For each accent, 10 utterances (5F/5M) were generated from three systems: CosyVoice2 (standard text), ParlerTTS (standard text), and CLARITY. Native speakers of each accent (fluent in English), rated corresponding samples (e.g., Chinese listeners for Chinese-accented English [^10] The setting assumes that Chinese listeners are more sensitive to Chinese-accented English speech.). There were 4F/4M listeners for CN/SG and 2F/2M for IN/GB. Samples were evaluated on a 1–5 scale for naturalness, accent accuracy, age, gender consistency, and overall instruction match.
@@
Fig. 3: RAAP accuracy (%) for gender, predicted accent, age attribute (left) and LLM-as-judge scores for various texts (right).
@@
Fig. 4: Accent accuracy (%) (left) and NISQA (right) for the baselines and proposed CLARITY with GPT adapted text.
@@
Table 1: Ablation study: RAAP with text similarity, accent confidence, and text adaptations. ‘Max’ means adapted text selection described in Eqn. 4.
Text Sim. Accent Score Text Apt. ACC (%) NISQA
@@
Fig. 5: Subjective Evaluation Results.
3.2 Experimental Results and Bias Analysis
Objective Evaluation:

### v2

@@ Abstract
Instruction-guided text-to-speech (TTS) research has reached a maturity level where excellent speech generation quality is possible on demand, yet two coupled biases persist in reducing perceived quality: accent bias, where models default towards dominant phonetic patterns, and linguistic bias, a misalignment in dialect-specific lexical or cultural information. These biases are interdependent and authentic accent generation requires both accent fidelity and correctly localized text. We present CLARITY (Contextual Linguistic Adaptation and Retrieval for Inclusive TTS sYnthesis), a backbone-agnostic framework to address both biases through dual-signal optimization. Firstly, we apply contextual linguistic adaptation to localize input text to align with the target dialect. Secondly, we propose retrieval-augmented accent prompting (RAAP) to ensure accent-consistent speech prompts. We evaluate CLARITY on twelve varieties of English accent via both subjective and objective analysis. Results clearly indicate that CLARITY improves accent accuracy and fairness, ensuring higher perceptual quality output [^1] Code and audio samples are available at https://github.com/ICT-SIT/CLARITY.
I INTRODUCTION
@@
Instruction-guided TTS systems [5, 11, 19, 17] taking two human inputs of (a) transcript text and (b) a speaker descriptions of desired output speech characteristics, have recently demonstrated promising performance. However, when prompted with natural language instructions, such TTS systems are susceptible to various biases. While gender bias has been examined by prior work [8], to the best of our knowledge, no study has jointly addressed the interrelated accent and linguistic bias problem. These biases are related, but their origins differ – linguistic bias is typically implicit in user input whereas accent bias stems from the TTS generation system. These two biases reduce the authenticity of generated speech primarily by misaligning what is being said with how it is spoken [1]. An example, illustrated in Fig. 1, might be to instruct the TTS to speak in a British Accent but supplying it with American English oriented text [^2] While users could supply better aligned TTS target text in this simple example, for less familiar language mixtures they may not know what is the appropriate text that aligns with the requested accent, and indeed this can be context dependent.. This example produces a confusing output of “British sounding American English”.
@@
Figure 1: The linguistic bias (prompt-source text) and accent bias (system-source acoustics) reduce both naturalness and perceived authenticity.
@@
Linguistic Bias – in the context of TTS, this bias is generally not an issue of model failure but a human-centric problem rooted in the diversity of users’ linguistic knowledge, shaped by education and culture [2]. It results in users supplying standardized or misaligned input text, even when dialectal diversity is intended. Recent studies confirm this tendency: for instance, ChatGPT frequently defaults to “standard” English when prompted with non-standard dialect text [7].
@@
Accent Bias – arises from training data imbalances, where accents well represented in the training set (e.g., American English) dominate, while less common ones (e.g. Malaysian English) are under-represented. As a result, models tend to generalise to mainstream accents even when instructed otherwise, making authentic accent generation difficult. Recent research confirms performance disparities across accents in synthetic speech. For example, Michel et al. [13] found significant quality gaps when cloning voices with different English accents, potentially reinforcing accent-based discrimination in AI speech. Broader evaluations like AudioTrust [10] revealed that accent is a stronger source of unfair bias in audio models than age or gender. Current English language TTS models generally default to American-accented speech profiles.
@@
CLARITY: To jointly mitigate the two interrelated biases, we propose Contextual Linguistic Adaptation and Retrieval for Inclusive TTS sYnthesis (CLARITY). This is a backbone-agnostic framework to enhance zero-shot TTS by providing richer guidance for both linguistic content and accent style. Part of its novelty lies in the joint optimization that mitigate the two biases through: (a) an LLM-guided text adaptation module to localize input to the target dialect, and (b) a metadata-driven retrieval mechanism to supply dialect-consistent accent prompts. CLARITY leverages large language models (LLMs) for on-the-fly text modification, via their knowledge of dialects, culture and slang, to enhance authenticity. The second innovation is to pair this with retrieval-augmented accent prompting, in to achieve consistent dialect-aware accent synthesis. To our knowledge, this is the first framework to jointly address user-side linguistic bias and system-side accent bias. Objective and subjective evaluations presented below demonstrate that these techniques allow CLARITY to improve fairness, accent fidelity, and overall authenticity.
II The Proposed framework
@@
As noted above, CLARITY jointly optimizes two signals: (a) an adapted text transcript $x$ and (b) dialect-consistent accent prompting $s$, as shown in Fig. 2. We cast this as a two-signal optimization problem that aims to maximize both linguistic and accent fidelity under the synthesis constraint of a backbone TTS model. Formally, the objective is defined as
@@
where $J_{\text{LLM}}(x^{*},m)$ denotes the judgment score assigned by an LLM-as-a-judge evaluating whether the adapted text $x^{*}$ aligns with the requested metadata $m$, and $C(s^{*},m)$ is the accent recognition confidence score that the retrieved prompt $s^{*}$ matches the target accent attributes. The synthesis function $g_{\text{TTS}}$ represents any zero-shot backbone TTS model. It is agnostic to TTS systems, and we will evaluate CLARITY in Section III with several alternatives.
II-A LLM-Guided Instruction Parsing
Let $u\in\mathcal{U}$ be a free-form user instruction in natural language, and $m\in\mathcal{M}$ the structured metadata schema, $m=(\text{accent}\in\mathcal{A},;\text{gender}\in{M,F},;\text{age}\in\mathbb{N},\ldots)$, where $\mathcal{A}$ is the set of supported accents. The challenge is that $u$ may encode attributes explicitly, vaguely, or implicitly; for example, instead of prescriptive instructions about age, gender and precise accent, it may be left open to interpretation such as u=u=Read this like a local in Singapore” (a multilingual, multiracial and multi-accent locale).
We first parse $u$ into structured metadata $m$,
@@ II-A LLM-Guided Instruction Parsing
where parsing function $f_{\text{parse}}$ is realized by prompting an LLM with schema-specific instructions. $\theta_{\text{LLM}}$ are its parameters. For each attribute $m_{j}$ (e.g., accent, gender, age), the LLM estimates a posterior distribution, $P(m_{j}\mid u;\theta_{\text{LLM}}),\quad m_{j}\in\mathcal{V}_{j},$ where $\mathcal{V}_{j}$ denotes the vocabulary/domain of attribute $j$. The predicted value is chosen as $m_{j}=\arg\max_{v\in\mathcal{V}_{j}}P(m_{j}=v\mid u;\theta_{\text{LLM}}).$ In cases where $u$ provides insufficient information to infer slot $m_{j}$, the system falls back to default $\pi_{j}$.
II-B Contextual Linguistic Text Adaptation
Let $x$ denote the input text provided by the user, linguistic bias on the user side arises because $x$ is often drawn from the distribution of the user’s learned or assumed dialect $\mathcal{D}_{u}$, which may differ from the target dialect $\mathcal{D}_{t}$. We formalize text adaptation as a conditional function as follows:
@@ II-B Contextual Linguistic Text Adaptation
To ensure that the adapted text $x^{\prime}$ is linguistically faithful to the target dialect $\mathcal{D}_{t}$, we first generate a set of candidate adaptations $\{x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{K}\}$ using different LLMs. For example, these might be GPT and LLaMA. We then compare the original standard text input $x_{\text{std}}$ with the LLM-adapted candidates, e.g. $k\in\mathcal{K}=\{\text{GPT},\text{LLaMA}\}$. An LLM-as-a-judge assigns each candidate a score $J_{\text{LLM}}(x,m)\in[0,10],$ which reflects how well $x$ aligns with metadata $m$, focusing on the accent attribute. The final text is then chosen as the candidate with the highest judgment score:
@@
Figure 2: The CLARITY framework for mitigating user-side linguistic bias (2.2) and system-side accent bias (2.3) to generate more authentic output (2.4) from user input (2.1).
II-C Retrieval-Augmented Accent Prompting (RAAP)
As discussed above, system-side accent bias arises when a TTS backbone trained on imbalanced data defaults to majority accents $\mathcal{A}_{\text{dom}}$. To counter this, we introduce Retrieval-Augmented Accent Prompting (RAAP), a RAG-like process which selects accent-consistent prompts from a curated pool $\mathcal{P}=\{(a_{i},z_{i},s_{i})\}$ containing metadata, transcript, and speech. Given target metadata $m$, we filter $\mathcal{P}$ by $a_{i}\approx m$ to form $\mathcal{P}_{m}$, then score each candidate $s_{i}$ using an ECAPA-TDNN [4] for accent confidence $C(s_{i},m)\in[0,1]$, and select those with the highest scores to guide accented speech generation. Mathematically this means,
@@ II-C Retrieval-Augmented Accent Prompting (RAAP)
To further align the text content of the prompt speech with the user’s text input, RAAP calculates the text similarity between the prompt speech candidates and the user-provided (standard) text input using $\cos(\phi(z_{i}),\phi(x))$, where $\phi(\cdot)$ is a TF–IDF [15] text embedding representation and $\cos(\cdot,\cdot)$ denotes cosine similarity. Specifically, we define the effective ranking score as,
@@
This ensures consistency in the meaning of the text, while allowing variation in word choice and sequence to better match the metadata.
II-D Backbone-agnostic Guided Synthesis
The final stage of CLARITY aligns the two signals: the adapted text $x$ and the accent-consistent prompt $s$, for guided synthesis. Let $g_{\text{TTS}}$ be the backbone TTS with parameters $\theta_{\text{TTS}}$. As a backbone-agnostic framework, CLARITY can use any instruction-guided, zero-shot TTS model. Synthesis is defined as,
@@ II-D Backbone-agnostic Guided Synthesis
III Experiments
III-A Experimental Settings
@@
Accent Pool Construction: We evaluated our method on twelve English accents. The accent pool is composed of accents from two datasets. Ten accents (Canadian (CA), Chinese (CN), Spanish (ES), British (GB), Indian (IN), Japanese (JP), Korean (KR), Portuguese (PT), Russian (RU), and American (US)) are from the AESRC dataset (52,614 utterances from 528 speakers) [16], each contributing 44–92 speakers and 4.2k–7k utterances, balanced by gender and aged 15–69. For our experiments, we selected a subset of 52,614 utterances, balanced by gender and covering a broad age range. For each speaker, up to 100 longest utterances were retained to ensure consistent recording quality and sufficient duration per sample. The remaining two accents are drawn from the SEAME dataset [12] of 40 Malaysian speakers (4,358 utterances, aged 20 to 33) and 114 Singaporean speakers (17,750 utterances, aged 18 to 24), with code-switching between English and Mandarin. To construct our subset, we extracted English-dominant utterances from both conversational and interview recordings, retaining only those labeled as English or code-switched (EN and CS) in the transcript metadata. Purely Mandarin (ZH) segments were excluded. Each utterance was required to contain at least 5 words, with a maximum of 100 utterances per speaker, prioritizing longer segments. After extraction, only speakers with English accent confidence scores greater than 0.9 were retained.
@@
Free-form Instructions Curation: For transcript generation, we used GPT-4333https://openai.com/index/gpt-4-research/ to formulate a set of standard sentences across four scenarios (restaurant, university, workplace, supermarket). Speaker instructions derived from accent-pool metadata (accent, age, gender) served both as prompts and ground-truth labels for retrieval. CLARITY adapted each instruction–text pair using GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini/ and LLaMA-3.1-8B [^5] https://huggingface.co/meta-llama/Llama-3.1-8B to produce accent-specific variant candidates. Prompts used for instruction curation is provided in the appendix.
@@
Baseline Systems: Two open-source SOTA prompt-based TTS systems are selected as baselines: ParlerTTS [9, 11] and CosyVoice2 [5]. Both take user instructions and standard text as input, but CosyVoice2 requires additional prompt speech even for instruction-guided functionality [^6] https://github.com/Render-AI/CosyVoice2 where a silent speech sample is provided.
@@
Object Evaluation Metrics: To quantify the accent level of the generated speech, we calculate accent accuracy using the ECAPA-TDNN accent recognition model [^7] https://huggingface.co/Jzuluaga/accent-id-commonaccent_ecapa, finetuned on the accent pool (Section III-A), as in RAAP. Speech quality is measured with NISQA [14] (on a 1–5 scale) using its pre-trained model [^8] https://github.com/gabrielmittag/NISQA. Bias/Fairness of the generated speech is evaluated with the Fairness Discrepancy Rate (FDR) [3, 6], which quantifies disparities in false alarm/reject rates across accents, with 1 indicating perfect fairness.
@@
Subject Evaluation Metrics: We conducted human listening tests on four of the accented English sample sets (CN, SG, IN, GB). For each accent, 10 utterances (5F++5M) were generated from three frameworks: CosyVoice2, ParlerTTS, and CLARITY (employing RAAP and linguistic text adaptation). Native speakers of each accent who are fluent in English, rated corresponding samples (e.g., Chinese listeners for Chinese-accented English [^9] The assumption is that Chinese listeners are more sensitive to Chinese-accented English speech and vice-versa.). The listening test volunteers comprised 4F++4M listeners for CN and for SG accented English while 2F++2M responded for each of IN and GB accented tests. Samples were evaluated on a 1–5 scale for naturalness, accent accuracy, age, gender consistency, and overall instruction match.
@@
Figure 3: RAAP accuracy for gender, predicted accent, age attribute (left) and LLM-as-judge scores (right).
@@
Figure 4: Accent accuracy (%) (left) and NISQA (right) for the baselines and proposed CLARITY with GPT adapted text. Higher scores indicate better results.
@@
TABLE I: Ablation study: RAAP with text similarity, accent confidence, and text adaptations. ‘Max’ means adapted text selection described in Eqn. 4.
[…]
@@
Figure 5: Subjective evaluation results for three instruction-guided TTS systems (CosyVoice2, ParlerTTS and the proposed CLARITY) on four dialects of English. The violin plots represent scores for attributes of quality (mean opinion score, MOS), accent, gender and age fidelity, as well as overall match.
III-B Objective Evaluation